OpenStudy (electrochika):

At time t equals or > 0, the acceleration of a particle moving on the x axis is a(t)=t+sint.?

OpenStudy (electrochika):

at t=0 the velocity of the particle is -2. for what value t will the velocity of the particle be zero?

OpenStudy (xishem):

To find a velocity equation from an acceleration equation, you need to differentiate the acceleration equation with respect to t: \[a'(t)=\cos(t)\]Now find the values for which this equation is equal to 0. \[0=\cos(t)\]\[t=\pi/2, 3\pi/2\]

OpenStudy (turingtest):

\[v(t)=\int a(t)dt=\frac1 2t^2-\cos t+C\]\[v(0)=-1-2=C=-3\]\[v(t)=\frac1 2t^2-\cos t-3=0\]solve for t

OpenStudy (anonymous):

think this might be backwards you are given acceleration and you want velocity

OpenStudy (turingtest):

Xishem has it backwards, yes

OpenStudy (anonymous):

what turing test said.

OpenStudy (xishem):

Yep.

OpenStudy (electrochika):

So t = - 2.917 and 2.917?

OpenStudy (turingtest):

yes, and I would imagine we want the positive answer here

OpenStudy (electrochika):

yeah we do

OpenStudy (turingtest):

oh yeah, positive only it says so in the problem